The efficiency of a heat engine if the temperature of the source is 100*c and sink is 27*c is nearly?
  1. 0.1
  2. 0.2
  3. 0.3
  4. 0.4

Grade:12

1 Answers

Arun
6 years ago
Dear student
Efficiency of heat engine = 1 - T2/T1
Hence efficiency = 1 -( 300/373 )
= 1 - 0.804
= O.196
Hence option B is correct.
Regards
Arun (askIITians forum expert)